OpenStudy (anonymous):

Please help, will medal -3x 4 _____ + ______ x2 – 9 2x – 6

OpenStudy (bibby):

get them in terms of a common denominator Start by factoring the left. recall that \(a^2-b^2=(a+b)(a-b)\)

OpenStudy (michele_laino):

please use these factorizations: \[\begin{gathered} {x^2} - 9 = \left( {x - 3} \right)\left( {x + 3} \right) \hfill \\ 2x - 6 = 2\left( {x - 3} \right) \hfill \\ \end{gathered} \]

OpenStudy (bibby):

also 4/2x-6=2/x-3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Then what @Michele_Laino

OpenStudy (michele_laino):

Using the factorization above, I get: \[\frac{{ - 3x}}{{\left( {x - 3} \right)\left( {x + 3} \right)}} + \frac{4}{{2\left( {x - 3} \right)}} = ...?\]

OpenStudy (michele_laino):

or, after a simplification: \[\frac{{ - 3x}}{{\left( {x - 3} \right)\left( {x + 3} \right)}} + \frac{2}{{\left( {x - 3} \right)}} = ...?\]

OpenStudy (anonymous):

ok yes I understand that much, but I don't know what to do with the denominator :/

OpenStudy (michele_laino):

what is the least common multiple?

OpenStudy (anonymous):

I've been very stressed and my brain just isn't understanding math lately, please remind me what a least common multiple is...

OpenStudy (michele_laino):

the least common multiple is: (x-3)(x+3)

OpenStudy (anonymous):

but why? I don't understand

OpenStudy (anonymous):

OH!! I just remembered thank you

OpenStudy (anonymous):

so does that mean you would have to multiple the right side by (x+3)?

OpenStudy (michele_laino):

since you have to pick common and uncommon factors with the highest exponent

OpenStudy (michele_laino):

it means that you have to multiply the numerator of the second fraction, by (x+3)

OpenStudy (anonymous):

so multiply both the numerator and denominator by (x+3)?

OpenStudy (michele_laino):

here is: \[\frac{{ - 3x}}{{\left( {x - 3} \right)\left( {x + 3} \right)}} + \frac{2}{{\left( {x - 3} \right)}} = \frac{{ - 3x + 2\left( {x + 3} \right)}}{{\left( {x - 3} \right)\left( {x + 3} \right)}}\]

OpenStudy (michele_laino):

note that the numerator of the first fraction is unchanged

OpenStudy (michele_laino):

so, what is: \[\frac{{ - 3x + 2\left( {x + 3} \right)}}{{\left( {x - 3} \right)\left( {x + 3} \right)}} = ...?\]

OpenStudy (anonymous):

the denominator is x^2-9 right?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

\[x^{2}-9\]

OpenStudy (anonymous):

:) and I'm not sure about the numerator...

OpenStudy (michele_laino):

you have to compute the multiplication first

OpenStudy (michele_laino):

what is 2(x+3)=...?

OpenStudy (anonymous):

2x+6

OpenStudy (anonymous):

so it would be -x+6?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (michele_laino):

Thank you!